Whispering Winds Corp. purchased a truck for $73,000. The company expected the truck to last four years or 110,000 miles, with an estimated residual value of $18,000 at the end of that time. During the second year the truck was driven 30,000 miles. Compute the depreciation for the second year under each of the methods below and place your answers in the blanks provided.

Units-of-activity $enter the depreciation for the second year under the Units-of-activity method in dollars

Double-declining-balance $enter the depreciation for the second year under the Double-declining-balance method in dollars
Answer #1

Units of activity

= (73,000-18,000)/110,000 * 30,000

15,000

Double declining balance

Rate = 200/Useful life"= 200/4 = 50%

First year = 73,000*50% =36,500

Second Year = (73,000-36,500)*50%

18,250